OpenStudy (anonymous):

Maura had a pizza delivered to her home. The total amount she paid was $17.53. Included in the total amount were the menu price of the pizza, 6% sales tax on the price of the pizza, and a $3.75 delivery fee. What was the menu price of the pizza? $

OpenStudy (anonymous):

k so what do we know first of all

OpenStudy (anonymous):

Okay first you need to determine what exactly was taxed. Since it say the pizza was taxed, you can deduct the delivery fee first.

OpenStudy (anonymous):

$ 17.53 - 3.75 = T where T is the price which was taxed. From here, T = 1.06*P where P is the price on the menu. The 1.06 comes from 100% of the original price plus the 6% of tax = 106% or 1.06

OpenStudy (anonymous):

well don't we need to add 17.53+6%

OpenStudy (anonymous):

nope u subtract $ 17.53 - 3.75 = T where T is the price which was taxed.

OpenStudy (anonymous):

wait sorry

OpenStudy (anonymous):

which is 13.78

OpenStudy (anonymous):

17.53 - 3.75 = 13.78 = T P = T/1.06 = 13 So the cost on the menu, before tax and delivery, is $13.

OpenStudy (anonymous):

so no add or subtract or thats the answer

OpenStudy (anonymous):

so $13

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

lets do another one k

OpenStudy (anonymous):

sure :)